OpenStudy (anonymous):

Simplify the following expression, express in terms of i sqrt(-48)

OpenStudy (anonymous):

\[\sqrt{-48}\]

OpenStudy (anonymous):

\[\sqrt{-48}=\sqrt{16\times 3\times (-1)}=\sqrt{16}\times \sqrt{3}\times \sqrt{-1}=4\sqrt{3}i\]

OpenStudy (anonymous):

the \(i\) is outside the radical

OpenStudy (anonymous):

thank you! I have been trying to figure out what number times itself is -48.......

OpenStudy (anonymous):

well you do not know an integer whose square is 48, so you just write in simplest radical form

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

algebra is so not my forte lol